1. The Earth is 8 light minutes from the sun and the speed of light is 3×10^8 m.s^−1. Let's assume that earth's orbit is exactly circular. Its mass is 5.972×10^24 kg.

Using the values given above (and your knowledge about time units), determine the magnitude of the gravitational force that the Sun exerts on the Earth.

The magnitude of the force is _____ newtons.

Format: if your answer is 2×10^3 N, enter 2*10^3 (and do think about significant figures).
1
Expert's answer
2020-03-23T10:42:43-0400

The gravitation force


"F=G\frac{mM}{r^2},"

"F=6.67\times 10^{-11}\frac{5.972\times 10^{24}\times 1.989\times 10^{30}}{(3\times 10^8\times 8\times 60)^2}"

"F=3.82\times 10^{22}\:\rm N."